An astronaut is being tested in a centrifuge. The centrifuge has a radius of 7.40 m and, in starting, rotates according to θ = 0.190t2, where t is in seconds and θ is in radians. When t = 3.60 s, what are the magnitudes of the astronaut's (a) angular velocity, (b) linear velocity, (c) tangential acceleration, and (d) radial acceleration?
Answer #1

Angular displacement of the astronaut is given by,

\theta(t) = 0.19t2

\omega = \frac{d \theta }{dt}

If we differentiate the equation of angular displacement with respect to time we will get the equation for angular velocity,

\omega(t) = 0.38t

\alpha = \frac{d\omega }{dt}

If we differentiate the equation of angular velocity with respect to time we will get the equation for angular acceleration,

\alpha(t) = 0.38

Radius of the centrifuge = R = 7.4 m

Angular velocity at t=3.6 sec = \omega1 = (0.38)(3.6) = 1.368 rad/s

Angular acceleration at t=3.6 sec = \alpha1 = 0.38 rad/s2

Linear velocity of the astronaut at t=3.6 sec = V1

V1 = \omega1R

V1 = (1.368)(7.4)

V1 = 10.12 m/s

Tangential acceleration of the astronaut at t=3.6 sec = at1

at1 = \alpha1R

at1 = (0.38)(7.4)

at1 = 2.812 m/s2

Radial acceleration of the astronaut at t=3.6 sec = ar1

ar1 = \omega12R

ar1 = (1.368)2(7.4)

ar1 = 13.85 m/s2

a) Magnitude of astronaut's angular velocity at t=3.6 sec = 1.368 rad/s

b) Magnitude of astronaut's linear velocity at t=3.6 sec = 10.12 m/s

c) Magnitude of astronaut's tangential acceleration at t=3.6 sec = 2.812 m/s2

d) Magnitude of astronaut's radial acceleration at t=3.6 sec = 13.85 m/s2